I want to confirm if I understood .
Problem states : Draw a right triangle with a base of 2 units and a height of 3 units. Make three “copies” so that the lengths of
the new sides are 50%, 300%, and 500% of the original.
Solution:
1st triangle
Base2×0.5= 1 new base will be 2+1=3
Height 3×0.5=1.5 new height 3+1.5=4.5
So new triangle base 3units and height 4.5 units
Second triangle:
2×3=6 so, 6+2=8
3×3=9 so, 3+9=12
New triangle base 8 height 12

Answer by josgarithmetic(39617)

To change your base of 2 and height of 3, to 300% of these, this means make the linear length measures 3 times bigger.

2*3=6 for the base measure
3*3=9 for the height measure


500% of the original?
new base is 2*5=10
new height is 3*5=15